# Language-based Valence and Arousal Expressions between the United States and China: a Cross-Cultural Examination

### v2

@@ 4 Discussion
One of the major findings in our study is that the American participants had stronger negativity bias and overall had higher arousal with higher positive and negative valence compared to Chinese participants. This is consistent with past findings on West-East distinction in emotional arousal: in Western or individualist culture, high-arousal emotions are valued and promoted more than low-arousal emotions, while in Eastern or collectivist culture, low-arousal emotions are valued more than high-arousal emotions Lim (2016). Even in traditional Asian medicine, there is an assumption that excessive emotional expression can be harmful and cause diseases, whether it is positive or negative emotions Lim et al. (2008). Our findings confirmed that Chinese users on Weibo express lower arousal levels for both negative and positive emotions.
@@ 5 Limitations
Even though Twitter and Weibo are comparable in usage Li et al. (2020); Guntuku et al. (2019c) and have not been shown to have significant differences in predicting individual states Gao et al. (2012), data from other platforms such as WeChat and RenRen in China and Facebook in the US have not been included in this study due to access constraints. Emojis are a significant contributor to affective expressions Li et al. (2019); however, we did not include them in this study due to differences in encodings while collecting the data making it infeasible for us to parse them accurately. Further, social media users are non-representative of the general population, and the participants in this study are non-random and convenient samples.
@@ 6 Ethics
All the figures and tables from Weibo are translated into English with Google Translate. In the Appendix, we show the figures with original Chinese text. Figure 6 shows the top 15 phrases for effect strength with valence and arousal. Figure 7 shows the top topics associated with valence and arousal.

### v4

@@ 4 Discussion
One of the major findings in our study is that the American participants had stronger negativity bias and overall had higher arousal with higher positive and negative valence compared to Chinese participants. This is consistent with past findings on West-East distinction in emotional arousal and aligns with Hofstede’s Individualism vs. Collectivism dimension: in Western or individualist culture, high-arousal emotions are valued and promoted more than low-arousal emotions, while in Eastern or collectivist culture, low-arousal emotions are valued more than high-arousal emotions Lim (2016); Hofstede (2011). This can be attributed to the fact that individualistic culture encourages expressive independence and the externalization of personal emotions while collectivist culture values social harmony and group cohesion. Even in traditional Asian medicine, there is an assumption that excessive emotional expression can be harmful and cause diseases, whether it is positive or negative emotions Lim et al. (2008). Our findings confirmed that Chinese users of Weibo express lower arousal levels for both negative and positive emotions.
@@ 6 Limitations
Platform Issue: Even though Twitter and Weibo are comparable in usage Li et al. (2020); Guntuku et al. (2019b) and have not been shown to have significant differences in predicting individual states Gao et al. (2012), data from other platforms such as WeChat and RenRen in China and Facebook in the US have not been included in this study due to access constraints. Emojis are a significant contributor to affective expressions Li et al. (2019); however, we did not include them in this study due to differences in encodings while collecting the data making it infeasible for us to parse them accurately. Further, social media users are non-representative of the general population, and the participants in this study are non-random and convenient samples.
Fine-grained Emotion: Our focus in this paper was to compare the expressions of valence and arousal across two different cultures building upon rich cross-cultural psychological studying the difference in valence and arousal Lim (2016); Kuppens et al. (2017a). Although considering fine-grained emotions could make the analysis multidimensional, it will make the results less reliable. Moreover, each of the fine-grained emotions could be represented in the valence and arousal circumplex Jefferies et al. (2008); Mohammad (2018b).
Subcultural Variance: We acknowledge the existence of subcultural variances, such as those among various ethnicities, provinces, and counties in China and the US. For example, minority students, including Tibetan and Mongolian, tend to experience more negative emotions and are less inclined to adopt emotion regulation strategies compared to Han students Lü and Wang (2012). Within the United States, European Americans show a greater motivation to engage in hedonic emotion regulation than their Asian American counterparts Miyamoto et al. (2014).
However, despite these nuances, the macro-cultural differences between East and West remain significant enough to warrant a comparative analysis Lu et al. (2001); Tsai et al. (2006b). Focusing on broader cultural variation, our investigation emphasizes the pronounced disparities between the US and Chinese cultural contexts by using social media posts. These disparities are substantial and provide a robust framework for comparative analysis. This macro-level perspective is not to negate the relevance of subcultural variances but to highlight the overarching patterns that emerge when contrasting Eastern and Western cultures by using two large countries that have a large variation in Hofstede’s cultural dimensions, for instance, as examples. By situating our work within this broader context, we aim to contribute to a more comprehensive understanding of how culture influences emotional expression. In the discussion, we will add the above language to acknowledge the complex variety of cultural diversity that exists within and across national borders.
Translation: Lexica need to be adapted to the cultures to measure psychological phenomenon accurately. We tried using Chinese valence-arousal words (CVAW, Lee et al. (2022)). However, we did not proceed further as the methods of building NRC-VAD (for English) and CVAW (for Chinese) lexica were different and could cause misalignment. We wanted to control for such differences by choosing a lexicon that has sufficient coverage while also being used in multiple prior works across both languages Wenjia et al. (2023); Mohammad (2016); Li et al. (2019); Chen et al. (2019); Das and Dutta (2021). Further, with over 20K entries, NRC-VAD is the largest manually created emotion lexicon that has translations across several languages.
@@ 7 Ethics
The data collection in our study were approved by the Institutional Review Board (IRB) of our institution.
@@
Appendix A Explanation of Conditional
In Table 1, we use conditional to represent the variance explained by the entire model, including both fixed and random factors.
The method we used in finding functional relationships between valence and arousal is a follow-up analysis using the same method used in Kuppens, Tuerlinckx, Russell, and Barrett (2013). Table 2 in the paper listed BIC, PostP, and the best for different studies and datasets, which are consistent with our result - relatively low scores. In our paper, we expand this work using social media data to see if a similar conclusion can hold across cultures with the functional relationship between valence and arousal.
As mentioned in Kuppens, Tuerlinckx, Russell, and Barrett (2013), low to moderate values of our functions are expected, not only because our analysis is based on noisy social media data, but also affective experiences of all combinations of valence and arousal can occur. For example, although less likely, the LDA result in our paper shows that positive valence low arousal states are represented by relaxing, and sleep.
Our goal in the analysis of functional relationships is not to train a model for predicting the arousal of a sentence using valence, but to explain the relationship between valence and arousal on average.
Appendix B Details of Valence-Arousal Relation Models
In this section, we show the full comparison between Model 1 to Model 6 in Figure 9. Among all, Model 6 fits the social media data best, shows negative offset and negative bias in the line of best fit. This is opposite to the previous findings where positive offsets were observed from lab-based measurements.
Appendix C Removal of NRC-VAD Lexicons
To give a multidimensional insight into culturally specific expressions of valence and arousal, we show Twitter part of Figure 3 without using words from NRC-VAD Lexicon in Figure 6. NRC-VAD contains 20,000 words, which covers most of the daily vocabulary. This figure, without lexicon words, consists of a lot of emojis, internet slang, and swear words.